# IDEA: Enhancing the Rule Learning Ability of Large Language Model Agent through Induction, Deduction, and Abduction

### v4

@@ 4 The IDEA Agent
The IDEA agent begins with an abductive action to generate an initial hypothesis, followed by a deduction step to create a new plan. Based on this plan, the agent interacts with the environment. Upon receiving feedback from the environment as a new observation, the agent may take an inductive action to refine the hypothesis or perform another interaction with the environment. Deductive action is invoked to adjust the plan every time the hypothesis changes. This reasoning loop continues until the puzzle is solved or a maximum number of steps is reached. After each step, the results are appended to the agent’s memory, including interaction outcomes and any modifications to the hypothesis or plan. We provide a simplified algorithm demonstrating how the IDEA agent operates in Algorithm 1. Specifically, at each step, we prompt the LLM to reflect on the information recorded in the IDEA agent’s components to make decisions and take actions. We employ the chain-of-thought (COT) reasoning (Wei et al., 2022) for all prompts. Detailed prompts for each type of action are available in Appendix A.4. More detailed implementation of the agent can be found in Appendix A.3.1. Similar to real-life scenarios, when agents solve tasks in RULEARN puzzles, they do not know the outcomes in advance. Consequently, it is challenging to decide when to refine or change their hypothesis and plans, as well as what interactive actions to take to gather pattern-revealing observations. A detailed example of the IDEA agent solving the Reactor Puzzle is provided in Figure 3.
@@ 4.1 Experiment Settings
To evaluate the effectiveness of IDEA, we respectively initialize it with three popular open-source LLMs, including Gemma-7B (Team et al., 2024), Llama3-8B, and Llama3-70B (Dubey et al., 2024), and two closed-source LLMs, GPT-3.5-Turbo (Dubey et al., 2024) and GPT-4o  [^1] https://openai.com/index/gpt-4o-system-card/. We compare the IDEA agent against the following two variants:
  • Baseline Agent: The Baseline agent lacks the reasoning loop of abduction, deduction, and induction and does not generate hypotheses or plans. At each step, it selects an interactive action solely based on its current memories and the goal.
  • Oracle-rule Agent: Even if the agent could successfully learn the correct rule, applying the learned rule to solve the puzzle is non-trivial. The Oracle-rule agent serves as a special control group to establish the Oracle performance with the ground-truth rule provided at the beginning. Specifically: 1) For the Function Operator puzzles, agents are given the exact forms of the functions. Their task is to derive the values of the coefficients. 2) For the Escape Room puzzles, agents are provided with how the password is constructed from the objects in the room. Their task is to derive the password using the provided rule. 3) For the Reactor puzzles, the reaction rule is given to the agents in natural language accompanied with examples. The agents only need to synthesize the target strings using the provided rules.
Each variant is evaluated on all three puzzle types (totaling 60 puzzles), with five trials per puzzle, resulting in 300 tests per variant. In our initial experiments, we found that using a higher temperature enhanced the agent’s creativity, allowing it to generate more diverse and higher-quality hypotheses that aided in discovering the underlying rules. Therefore, we set the temperature to $1.0$. We apply the COT for the prompting of each step for all agents. Additionally, since we observed that the success rate does not improve after 15 interactive steps, we set the maximum interaction step count to 15. The agent is considered to have failed a puzzle if it doesn’t solve it within 15 interactive steps.
@@ 4.2 Human participants
To compare the behavior and performance of humans with that of LLMs in the abduction, deduction, and induction reasoning processes, we recruited 50 participants and randomly selected 30 puzzles (10 from each of the three puzzle types), assigning each participant three puzzles—one from each type. Each puzzle was attempted by five different individuals, ensuring no prior knowledge of the underlying rules and maintaining an unbiased sample across puzzle types. Details about Institutional Review Board (IRB) approval and participant recruitment are provided in §7. The human participants are instructed to operate under the same reasoning procedure as IDEA. Failure to solve a puzzle within 15 steps resulted in marking the attempt as unsuccessful, ensuring a fair comparison.
@@ 4.3 Main Results
We calculated the Wilson Confidence Interval (CI) (Brown et al., 2001) for puzzle solving success rate across different variants. The detailed results are displayed in Table 3.
Table 3: Puzzle Success Rate (95% CI). The table displays the success rate for each setting, with each pair of numbers in parentheses representing the lower and upper bounds of the 95% confidence interval for the success rate. The human results are based on a randomly selected 50% subsample of the puzzles. We can see that across all LLMs, IDEA achieves consistent significant improvements except for the Reactor puzzles. We use Boldface to highlight the performance comparisons between Baseline and IDEA Agents with GPT-4o.
@@
Oracle-rule Agent Gemma-7B (0.71, 3.84) (0.0, 3.7) (1.57, 9.84) (0.18, 5.45)
Llama3-8B (7.09, 13.92) (2.78, 12.48) (12.51, 27.78) (2.15, 11.18)
Llama3-70B (33.33, 44.29) (28.18, 46.78) (48.21, 67.2) (14.17, 29.98)
GPT-3.5-Turbo (3.83, 9.28) (0.55, 7.0) (4.11, 15.0) (4.11, 15.0)
GPT-4o (47.35, 58.57) (35.61, 54.76) (72.22, 87.49) (24.56, 42.69)
Baseline Agent Gemma-7B (0.34, 2.9) (0.0, 3.7) (0.18, 5.45) (0.55, 7.0)
Llama3-8B (2.55, 7.27) (0.55, 7.0) (4.81, 16.23) (0.55, 7.0)
Llama3-70B (18.6, 28.09) (22.78, 40.63) (12.51, 27.78) (12.51, 27.78)
GPT-3.5-Turbo (3.57, 8.89) (5.52, 17.44) (1.03, 8.45) (1.57, 9.84)
GPT-4o (19.21, 28.79) (18.6, 28.09) (14.17, 29.98) (14.17, 29.98)
IDEA Agent (Ours) Gemma-7B (0.06, 1.86) (0.18, 5.45) (0.0, 3.7) (0.0, 3.7)
Llama3-8B (1.36, 5.17) (0.18, 5.45) (3.43, 13.75) (0.0, 3.7)
Llama3-70B (25.41, 35.76) (34.67, 53.77) (21.89, 39.58) (10.89, 25.55)
GPT-3.5-Turbo (10.53, 18.38) (22.78, 40.63) (2.78, 12.48) (2.15, 11.18)
GPT-4o (30.14, 40.9) (30.94, 49.8) (32.8, 51.79) (16.69, 33.23)
Human (55.38, 70.62) (52.15, 77.56) (42.31, 68.84) (54.19, 79.24)
For the Oracle-rule agent, GPT-4o exhibits superior deductive capabilities, outperforming other models with an approximate 50% success rate across all three puzzle types. Specifically, in the Escape Room puzzles, agents achieve up to an 80% success rate by simply following the provided rule to count the number of paintings of specific colors and types. However, in the Function Operator and Reactor puzzles, merely knowing the rule is not sufficient for success, applying the rule to solve the puzzle is more challenging. Agents must also devise effective plans to apply these rules to reach a solution. This additional planning requirement results in a comparatively lower success rate in these puzzles.
@@
For the Baseline agent, where they are not provided with the underlying rules and solely rely on historical observations to make interactive actions, GPT-4o continues to lead in performance. Notably, Llama3-70B narrows the performance gap, especially in the Function Operator puzzles. Across models, the success rates drop by about half compared to the Oracle-rule agent. This significant decrease highlights the challenge of rule learning and indicates that current LLM agents struggle to learn rules in unfamiliar environments without explicit guidance.
@@
The IDEA Agent significantly boosts success rates. Our proposed IDEA framework leads to approximate 10% increases in success rates for Llama3-70B, GPT-3.5-Turbo, and GPT-4o compared to the Baseline agent. This improvement demonstrates that incorporating a reasoning loop of abduction, deduction, and induction substantially enhances the LLM rule-learning performance in unfamiliar environments. IDEA enables the LLMs to generate hypotheses, plan actions, and refine their understanding based on new observations, which is crucial for rule learning. However, smaller models like Llama3-8B and Gemma-7B experience a slight decline in performance when applying IDEA. This decline is likely due to the increased complexity and context management demands introduced by the IDEA framework, which pose challenges for small models with limited capacity.
@@
LLM agents still fall far behind humans. In the Escape Room puzzle, where the primary challenge is to discover the rule, the Oracle-rule agents excel because once the rule is identified, applying the rule is simple. However, in other types of puzzles, human participants significantly outperform all LLM agents, including the Oracle-rule agents, even without knowing the rules beforehand. This notable disparity highlights that LLM agents lag substantially behind humans, not only in learning new rules but also in effectively applying known rules to solve problems.
@@ 4.4 Analysis
When focusing on human participants, we observe that they solve fewer puzzles in the initial steps compared to the LLM agents. However, as interactions continue, the number of puzzles solved by humans increases rapidly, eventually surpassing that of all LLM agents. In contrast, LLM agents solved 88.76% of the puzzles within the first 10 steps. Beyond this point, additional interactions contribute less to their success rate. This pattern suggests that humans have a superior ability to learn continuously from interactive environments, effectively improving their performance over time. If we did not limit the puzzles to 15 steps, we anticipate that the success rate of human participants would be even higher. We will explore this observation further in Section 5.
IDEA reduces repetitive actions. LLM agents frequently repeat previous actions instead of exploring new ones. This behavior is highly inefficient in our controlled puzzle environments, where each interaction yields deterministic results, and repeating the same action generally does not provide new information. An exception is the Escape Room puzzle, where entering the same password could result in different outcomes after the rule changes following three failed attempts.
We calculate the average number of repeated actions performed while solving each puzzle, with detailed statistics presented in Table LABEL:tab:repeat_count in Appendix A.2. We observe that most LLMs commonly repeat actions in the Baseline agent. Although models like GPT-4o and Llama3-70B exhibit this issue to a lesser degree, even these advanced models sometimes repeat actions that fail to provide further insights. The IDEA agent effectively reduces this tendency by explicitly generating plans during the deduction phase. By outlining a clear plan, the IDEA agent can better assess whether the current observations are sufficient or if further specific evidence is needed to reveal the underlying rule. For example, in the Escape Room puzzle, the IDEA agent avoids unnecessary attempts at entering passwords when the evidence gathered is sufficient to determine the correct code (see Figure 11 in Appendix A.4).
@@ 5 Human evaluation
Abduction stage. During the abduction stage, agents formulate an initial hypothesis based on initial observations. Since some puzzles are simple, agents can potentially guess the ground truth rule at this stage. Figure 5(a) shows that LLMs such as GPT-4o have approximately a 30% success rate in correctly identifying the rule during abduction. Notably, humans exhibit the lowest success rate in this stage. We believe this is because humans often lack the patience to thoroughly review all textual information initially and may overlook certain details. In contrast, LLMs meticulously process every word of the prompt and leverage their comprehensive pre-training data to generate plausible hypotheses. Additionally, humans tend not to formalize hypotheses when operating under uncertainty.
Deduction stage. After establishing an initial hypothesis—or each time the agent refines a hypothesis—the agent derives a plan to either validate this hypothesis or use it to solve the puzzle. As shown in Figure 5(b), humans generally outperform LLMs in creating high-quality plans. However, in the Function Operator puzzles, Llama3-70B and GPT-4o perform better than humans, suggesting that these models may have superior mathematical intuition compared to our human participants, who were undergraduate students from computer science-related majors.
Induction stage. Figure 5(c) illustrates the proportion of inductions where the refined hypothesis improved over the previous one, which we define as the effective induction rate. Humans outperform LLMs by a significant margin in this process, with 40% of their refined hypotheses being better than the previous ones. Surprisingly, GPT-4o performs worse than Llama3-70B and GPT-3.5-Turbo. We found that GPT-4o tends to hallucinate, especially in Escape Room puzzles, which negatively impacts its overall effective induction rate. In contrast, Llama3-70B almost never engages in induction activities within Reactor puzzles, as it fails to recognize when its hypothesized rules contradict its observations (see Appendix A.4.4). Despite these issues, both Llama3-70B and GPT-4o exhibit performance closer to humans during the induction process compared to other models in Function Operator puzzles. This relative success is believed to be due to the extensive training of LLMs with mathematical content, which ensures stable and superior performance in these types of puzzles. In contrast, Escape Room and Reactor puzzles, which are less likely to appear extensively in training datasets, see poorer and more variable performance from the models.
Average interactions needed for effective induction. Figure 5(d) shows that humans require fewer interactions—approximately four on average—to effectively refine their hypotheses, compared to LLMs. While LLMs can process initial information thoroughly and generate plausible hypotheses, they face challenges in refining these hypotheses based on new observations during interaction with the environment (see Figure 7 in Appendix A.1). This limitation suggests that LLMs may struggle to learn from new observations and incorporate feedback to continuously improve their hypotheses and problem-solving strategies. This gap may become more pronounced when agents are faced with larger action spaces and environments governed by more complex rules.
@@ 6 Conclusion
In this work, we introduce RULEARN, a benchmark designed to evaluate LLM agents’ rule-learning abilities in interactive environments. We propose IDEA, an agent framework that mimics human reasoning through abduction, deduction, and induction. Comprehensive experiments involving five prominent LLMs and human participants reveal that while IDEA significantly improves the rule learning ability of LLM agents, there is still a large gap between LLM and humans particularly in refining hypotheses and adapting strategies. Despite these advancements from the IDEA framework, LLMs still face challenges in generating valid hypotheses and avoiding repetitive actions in complex scenarios. Our findings underscore the need for further development of LLMs that can emulate human cognitive processes more effectively in explorations of novel environments. RULEARN provides a foundational resource for future research aimed at closing these gaps.

### v5

@@ 4 The IDEA Agent
The IDEA agent begins with an abductive action to generate an initial hypothesis, followed by a deduction step to create a new plan. Based on this plan, the agent interacts with the environment. Upon receiving feedback from the environment as a new observation, the agent may take an inductive action to refine the hypothesis or perform another interaction with the environment. Deductive action is invoked to adjust the plan every time the hypothesis changes. This reasoning loop continues until the puzzle is solved or a maximum number of steps is reached. After each step, the results are appended to the agent’s memory, including interaction outcomes and any modifications to the hypothesis or plan. We provide a simplified algorithm demonstrating how the IDEA agent operates in Algorithm 1. Specifically, at each step, we prompt the LLM to reflect on the information recorded in the IDEA agent’s components to make decisions and take actions. We employ the chain-of-thought (CoT) reasoning (Wei et al., 2022) for all prompts. Detailed prompts for each type of action are available in Appendix A.5.
More detailed implementation of the agent can be found in Appendix A.3.1. Similar to real-life scenarios, when agents solve tasks in RULEARN puzzles, they do not know the outcomes in advance. Consequently, it is challenging to decide when to refine or change their hypothesis and plans, as well as what interactive actions to take to gather pattern-revealing observations. A detailed example of the IDEA agent solving the Reactor Puzzle is provided in Figure 5 in Appendix.
5 Experiment Results
@@ 5.1 Experiment Settings
To evaluate the effectiveness of IDEA, we respectively initialize it with three popular open-source LLMs, including Gemma-7B (Team et al., 2024), Llama3-8B, and Llama3-70B (Dubey et al., 2024), and two closed-source LLMs, GPT-3.5-Turbo (Dubey et al., 2024) and GPT-4o [^2] https://openai.com/index/gpt-4o-system-card/. We compare IDEA against the following two variants:
  • ReAct Agent (Baseline): We choose ReAct (Yao et al., 2023b) as our baseline. The ReAct agent does not incorporate the full reasoning loop of abduction, deduction, and induction, nor does it generate explicit hypotheses or plans. Instead, at each step, it reasons over its current memories and the goal and selects an interactive action accordingly.
  • Oracle-rule Agent: Even if the agent could successfully learn the correct rule, applying the learned rule to solve the puzzle is non-trivial. The Oracle-rule agent serves as a control group to establish the Oracle performance with the ground-truth rule provided. Specifically: 1) For the Function Operator puzzles, agents are given the exact forms of the functions. Their task is to derive the values of the coefficients. 2) For the Escape Room puzzles, agents are provided with how the password is constructed from the objects. Their task is to derive the password using the provided rule. 3) For the Reactor puzzles, the reaction rule is given to the agents in natural language accompanied by examples. The agents only need to synthesize the target strings.
Each variant is evaluated across all three puzzle types. We set the temperature for LLMs to 0, based on observations that models like GPT-4o perform better at lower temperatures. The prompts used for the agents are detailed in Appendix A.5. Additionally, since the success rate does not improve after 15 interactive steps for LLMs, we capped the maximum interaction step count at 15. An agent is considered to have failed a puzzle if it does not solve it within these 15 steps. Details on the computational budget are available in Appendix A.4.
@@ 5.2 Human participants
To compare human and LLM performance in abduction, deduction, and induction reasoning, we recruited 50 participants and assigned each three randomly selected puzzles (10 from each reasoning type). Each puzzle was attempted by five different participants, with no prior exposure to the rules. Participants followed the same reasoning procedure outlined in IDEA, which mirrors their natural problem-solving methods and does not bias their responses. They documented their reasoning processes, enhancing transparency and facilitating clearer comparisons with LLMs. Further details on IRB approval and participant recruitment are in §9. Attempts failing to solve a puzzle within 15 steps were marked as unsuccessful, ensuring fair comparisons. A sample user interface is shown in Figure 18 in Appendix A.6.
@@ 5.3 Main Results
We calculated average puzzle solving success rate across different variants. The detailed results are displayed in Table 3.
Table 3: Puzzle Success Rate. The success rates for each setting. Across all LLMs, IDEA achieves consistently significant improvements, except for Gemma-7B in the Reactor puzzles. We use boldface to highlight performance comparisons between the Baseline and IDEA agents with GPT-4o.
@@
Oracle-rule Agent Gemma-7B $1.67$ $0.0$ $5.0$ $0.0$
Llama3-8B $5.67$ $1.0$ $14.0$ $2.0$
Llama3-70B $32.67$ $33.0$ $48.0$ $17.0$
GPT-3.5-Turbo $6.33$ $7.0$ $11.0$ $1.0$
GPT-4o $66.0$ $77.0$ $91.0$ $30.0$
ReAct Agent (Baseline) Gemma-7B $0.33$ $0.0$ $0.0$ $1.0$
Llama3-8B $1.67$ $0.0$ $5.0$ $0.0$
Llama3-70B $19.67$ $33.0$ $17.0$ $9.0$
GPT-3.5-Turbo $5.33$ $13.0$ $3.0$ $0.0$
GPT-4o $\mathbf{43.33}$ $\mathbf{62.0}$ $\mathbf{45.0}$ $\mathbf{23.0}$
IDEA Agent (Ours) Gemma-7B $0.33$ $0.0$ $1.0$ $0.0$
Llama3-8B $4.33$ $7.0$ $5.0$ $1.0$
Llama3-70B $29.0$ $41.0$ $35.0$ $11.0$
GPT-3.5-Turbo $7.33$ $18.0$ $3.0$ $1.0$
GPT-4o $\mathbf{50.33}$ $\mathbf{73.0}$ $\mathbf{51.0}$ $\mathbf{27.0}$
Human $63.33$ $66.0$ $56.0$ $68.0$
@@
For the Oracle-rule agent, in the Escape Room puzzles, agents achieve up to an 89% success rate by simply following the provided rule. However, in other puzzles, merely knowing the rule is not sufficient for success; applying the rule to solve the puzzle is more challenging. The Baseline agent is not provided with the underlying rules and solely relies on historical observations to make interactive actions. Across models, the success rates drop by about half compared to the Oracle-rule agent. This significant decrease highlights the challenge of rule learning and indicates that current LLM agents struggle to learn rules in unfamiliar environments without explicit guidance.
@@
IDEA significantly boosts success rates. Our proposed IDEA framework leads to approximate 10% increases in success rates for Llama3-70B, GPT-3.5-Turbo, and GPT-4o compared to the Baseline agent. This improvement demonstrates that incorporating a reasoning loop of abduction, deduction, and induction substantially enhances the LLM rule-learning performance in unfamiliar environments. IDEA enables the LLMs to generate hypotheses, plan actions, and refine their understanding based on new observations, which is crucial for rule learning. However, smaller models like Llama3-8B and Gemma-7B do not perform better when applying IDEA. This is likely due to the increased complexity and context management demands introduced by the IDEA framework, which pose challenges for small models with limited capacity.
@@
LLM agents still fall far behind humans. In the Escape Room puzzle, where the primary challenge is to discover the rule, the Oracle-rule agents excel because once the rule is identified, applying the rule is simple. However, in other types of puzzles, human participants significantly outperform all LLM agents, including the Oracle-rule agents, even without knowing the rules beforehand.
@@ 5.4 Analysis
When focusing on human participants, we observe that they solve fewer puzzles in the initial steps. However, as interactions continue, the number of puzzles solved by humans increases rapidly, eventually surpassing that of all LLM agents. In contrast, LLM agents solved 88.76% of the puzzles within the first 10 steps. Beyond this point, additional interactions contribute less to their success rate. This pattern suggests that humans have a superior ability to learn continuously from interactive environments, effectively improving their performance over time. If we did not limit the puzzles to 15 steps, we anticipate that the success rate of human participants would be even higher.
IDEA reduces repetitive actions. LLM agents frequently repeat previous actions instead of exploring new ones. This behavior is highly inefficient in our controlled puzzle environments, where each interaction yields deterministic results, and repeating the same action generally does not provide new information. We calculate the average number of repeated actions performed while solving each puzzle, with detailed statistics in Table LABEL:tab:repeat_count in Appendix A.2. We observe that most LLMs commonly repeat actions in the Baseline agent. The IDEA agent effectively reduces this tendency by explicitly generating plans during the deduction phase. By outlining a clear plan, the IDEA agent can better assess whether the current observations are sufficient or if further specific evidence is needed to reveal the underlying rule. For example, in the Escape Room puzzle, the IDEA agent avoids unnecessary attempts at entering passwords when the evidence gathered is sufficient to determine the correct code (see Figure 11 in Appendix A.5).
@@ 6 Human evaluation
Abduction stage. In this stage, agents formulate an initial hypothesis based on initial observations. Since some puzzles are simple, agents can potentially guess the ground truth rule at this stage. Figure 4(a) shows that LLMs such as GPT-4o have approximately a 30% success rate in correctly identifying the rule during abduction. Notably, humans exhibit lower success rate in this stage. We believe this is because humans often lack the patience to thoroughly review all textual information initially and may overlook certain details. In contrast, LLMs meticulously process every word of the prompt and leverage their comprehensive pre-training data to generate plausible hypotheses. Additionally, humans tend not to formalize hypotheses when operating under uncertainty.
Deduction stage. After establishing an initial hypothesis—or each time the agent refines a hypothesis—the agent derives a plan to either validate this hypothesis or use it to attempt the puzzle. As shown in Figure 4(b), humans generally outperform LLMs in creating high-quality plans.
Induction stage. Figure 4(c) illustrates the proportion of inductions where the refined hypothesis improved over the previous one, which we define as the effective induction rate. The induction stage is crucial in rule-learning since humans typically develop high-quality hypotheses through induction after experiments. However, LLM agents perform poorly at this stage. Humans significantly outperform LLMs, with 40% of their refined hypotheses being better than the previous ones. Surprisingly, GPT-4o performs worse than Llama3-70B and GPT-3.5-Turbo. We found that GPT-4o tends to hallucinate, especially in Escape Room puzzles, which negatively impacts its overall effective induction rate. In contrast, Llama3-70B almost never engages in induction within Reactor puzzles, as it fails to recognize when its hypothesized rules contradict its observations (see Appendix A.6.1).
Average interactions needed for effective induction. Figure 4(d) shows that humans require fewer interactions—approximately four on average—to effectively refine their hypotheses, compared to LLMs. While LLMs can process initial information thoroughly and generate plausible hypotheses, they face challenges in refining these hypotheses based on new observations during interaction with the environment (see Figure 7 in Appendix A.1). This limitation suggests that LLMs may struggle to learn from new observations and incorporate feedback to continuously improve their hypotheses and problem-solving strategies. This gap may become more pronounced when agents are faced with larger action spaces and more complex rules.
@@ 7 Conclusion
In this work, we introduce RULEARN, a benchmark to evaluate LLM agents’ rule-learning abilities in interactive environments. We propose IDEA, an agent framework that mimics human reasoning through abduction, deduction, and induction. Comprehensive experiments involving five prominent LLMs and human participants reveal that while IDEA significantly improves the rule learning ability of LLM agents, there is still a large gap between LLM and humans particularly in refining hypotheses and adapting strategies. Despite these advancements from the IDEA framework, LLMs still face challenges in generating valid hypotheses and avoiding repetitive actions in complex scenarios. Our findings underscore the need for further development of LLMs that can emulate human cognitive processes more effectively in explorations of novel environments. RULEARN provides a foundational resource for future research aimed at closing these gaps.
8 Limitations
While solving puzzles, the IDEA agent needs to manage long contexts. As exploration progresses and the agent encounters more observations, it must simultaneously process all observations. This requirement can limit its effectiveness in scenarios that involve lengthy contexts and complex rules, where extensive experimentation is needed to uncover these rules. By prioritizing and focusing on more critical observations, we can enhance the IDEA agent’s performance in managing long-context scenarios and in tackling challenging puzzles that require multiple steps to gather sufficient evidence.
@@ 9 Ethics Statement
10 Acknowledgements
During the final preparation of this manuscript, we utilized the GPT-4 language model provided by OpenAI to assist in identifying and correcting typographical and grammatical errors. The use of this tool was restricted solely to the polishing stage and did not influence the study’s conceptual framework, research methodology, data analysis, or conclusions. All substantive content and intellectual contributions remain those of the authors, and the AI assistance served only to ensure clarity and precision in the final written presentation.